OpenStudy (teller):

I need some help with this question not quite understanding how to do this .I suck at word problems in general . A family has two cars. The first car has a fuel efficiency of 15 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 1900 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (a_clan):

Okay. break the problem into smaller problems. 1)The first car has a fuel efficiency of 15 miles per gallon of gas. 2)the second has a fuel efficiency of 40 miles per gallon of gas. 3)the two cars went a combined total of 1900 miles. 4)total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars? From (1) in 1 gallon = 15 miles . Let it have consumed 'x' gallons in one week . Then total distance travelled in one week = '15x' (2) in 1 gallon = 40 miles .Let it have consumed 'y' gallons in one week . Then total distance travelled in one week = '40y' (3) Total = 1900 miles. Means 15x + 40y = 1900 .. equation (3) (4) Total gas consumed = 60 gallons. Then x + y = 60 ....equation (4) Now solve for equation (3) and (4)

OpenStudy (teller):

Would number three be 55 and for number for I'm not quiet sure...Mind helping me with another one ?

OpenStudy (teller):

four

OpenStudy (a_clan):

These are called simultaneous equations, which need to be solved together. If you don't know how to do them, you should learn this first.

OpenStudy (teller):

I actually can't learn them due to the fact my course I'm taking doesn't give lessons like a I don't know what it's called it;s like they give you questions like practice questions and you're supposed to solve them .I'm on the last part which is word problems and their kind of hard to understand math isn't my strongest subject .

OpenStudy (a_clan):

You won't be able to understand word problems unless you learn to solve the basic methodology.

OpenStudy (teller):

Mind helping me with another one then .I think I only have two left...

OpenStudy (a_clan):

post your questions. I will see what I could do.

OpenStudy (teller):

A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 75 pounds. The truck is transporting 55 large boxes and 60 small boxes. If the truck is carrying a total of 4200 pounds in boxes, how much does each type of box weigh?

OpenStudy (teller):

For me the first thing that comes to mind is adding like 75 to 4200 but i'm guessing that is wrong...

OpenStudy (a_clan):

This is the basic. You have to assume the unknowns as x,y,z and so on. For example, let the weight of large box be x and small box be y. Then , it is given x+y = 75. .......this is equation (1) Secondly, 55 large boxes weigh = 55x and 60 small boxes weigh = 60y Truck is carrying total 4200 lbs, then 55x + 60y = 4200 .....this is equation (2) Now , it remains to solve equation (1) and (2) These are two equations in two variable type problems.

OpenStudy (teller):

Alright so how would you solve them...

OpenStudy (a_clan):

refer this site for the fundamentals http://www.purplemath.com/modules/systlin5.htm

OpenStudy (teller):

Little confusing but don't know .That's why I come to the site to get help.But thanks I guess.